# Remarks on Dirac–Bergmann algorithm, Dirac’s conjecture and the extended Hamiltonian

### v2

@@ 5 Generalization to all canonical mechanical gauge systems
Hence, the right-hand sides of the constraints in the Lagrangian formalism become spurious. As we have seen, this can be interpreted as additional matter contributions in Yang–Mills theory (39) and in GR (67). We also find it curious that fixing a gauge in terms of the Lagrange multipliers $\lambda^{2}_{a}$ with $\lambda^{2}_{a}\neq 0$ leads to a modification of the original system in which the constraints are relaxed [22].
Returning to the case $f^{a}=f^{a}(q_{b})$, it is interesting to note that the Lagrangian constraints do not depend on the unresolved velocities, $\phi^{2}_{a}=\phi^{2}(q,\dot{q}_{A})$, and can therefore be brought to a form that depends only on the coordinates via a gauge transformation [^22] This is a consequence of the form of Castellani generator for such systems $G_{C}=\dot{\varepsilon}\ [p^{a}-f^{a}(q_{b})]+\varepsilon\ G_{1}$, where $G_{1}$ includes $\phi^{a}_{2}(q_{A},p^{B})$. We can set all resolved velocities to zero with the help of $G_{1}$., $\phi^{2}_{a}=\phi^{2}(q)$. This implies that they correspond to static solutions, or, in the case of field theories, to static configurations of the fields. They are lost in this incorrect interpretation.
@@
The argument generalizes to the case when the gauge “hits” arbitrarily many times. The gauge can hit more than twice in the sense that tertiary constraints $\phi^{a}_{3}$ may arise from the preservation of the secondary constraints $\phi^{a}_{2}$, followed by further constraints originating from the preservation of the tertiary ones, and so on, with all constraints being first-class [^23] An example of this is a system $L=\dot{x}\dot{z}+yz$ considered in [19].. The proof in such situations is analogous: one performs a procedure similar to (71)
$q_{i}=q^{\scriptscriptstyle E}_{i}-\lambda^{3}_{a}\{q^{\scriptscriptstyle E}_{i},\phi^{a}_{2}\},\qquad p^{i}=p_{\scriptscriptstyle E}^{i}-\lambda^{3}_{a}\{p_{\scriptscriptstyle E}^{i},\phi^{a}_{2}\}.$ (74)
@@
The tertiary constraints can be chosen as $\phi^{a}_{3}\equiv\{\phi^{a}_{2},H_{can}\}$. After performing this substitution and retaining terms only to first order in $\lambda^{3}_{a}$, one again arrives at $L^{(1)}=p^{i}\dot{q}_{i}-H_{E}$.
@@
Moreover, the above results extend to systems that possess first–class constraints together with second–class ones. Second–class constraints make it possible to impose restrictions on variables belonging to a single canonically conjugate pair simultaneously and to eliminate these variables from the equations. This elimination can be carried out simply by direct substitution into the total Hamiltonian [6], thereby reducing the system to one with only first–class constraints [^24] Second–class constraints also correspond to constrained modes. Eliminating the constraints in this manner does not remove the corresponding information they carry from the theory. They must still be accounted for, since they often correspond to observable quantities [17].. In practice, however, this procedure may turn out to be rather cumbersome.
@@
Thus, we have shown that all our arguments apply to all canonical mechanical gauge theories [^25] A trivial exception occurs in theories in which the gauge “hits” only once. The only possible first-class constraints in such theories are primary ones. In that case the total and extended Hamiltonians coincide, and Dirac’s conjecture is automatically satisfied. An example of this is a relativistic free particle.. Carrying out an analogous generalization to the case of field theories also presents no difficulty.
@@ 6 Conclusions
In this work, we have explored certain aspects of the Dirac–Bergmann approach to the Hamiltonian formulation of gauge theories that, in our view, do not receive the attention they deserve in the literature. We conclude that the question of gauge symmetry generators for degenerate Hamiltonian systems, and the treatment of such systems within the extended Hamiltonian framework, is more subtle than it may appear. Ignoring these nuances leads to a naïve interpretation where the canonical variables are assumed to play the same role as in the total Hamiltonian approach. In this case, the system appears to possess twice the actual gauge symmetry, since Dirac’s conjecture is fulfilled, with the constrained modes becoming spurious.
Although there is nothing inherently wrong with introducing additional gauge freedom into a theory, and even though such a naïve interpretation leaves the dynamical modes intact, it is essential to account for the constrained part of the physical sector, as it corresponds to non-dynamical observables. Constrained modes must therefore be expressed in terms of combinations that commute with the first–class constraints in the extended Hamiltonian approach. This actually implies that we must redefine the system’s observables. Interestingly, this redefinition can be interpreted as a version of the Stückelberg trick applied to variables that are second-class with respect to the set of all primary first-class constraints.

### v3

@@ 5 Generalization to all canonical mechanical gauge systems
As we have seen, this can be interpreted as additional matter contributions in Yang-Mills theory (39) and in GR (68). We also find it curious that fixing a gauge in terms of the Lagrange multipliers $\lambda^{2}_{a}$ with $\lambda^{2}_{a}\neq 0$ leads to a modification of the original system in which the constraints are relaxed [23].
Furthermore, to clarify how the system is modified upon transitioning to the extended Hamiltonian, it is useful to examine the corresponding modification of the Lagrangian. In the case of the total Hamiltonian, the Lagrangian is recovered from the first-order action by substituting the primary constraints and momentum definitions while identifying the unresolved velocities with Lagrange multipliers. A similar procedure can be applied to the extended Hamiltonian by treating the extended Hamiltonian system as equivalent to a total Hamiltonian system in an enlarged phase space $(q^{\scriptscriptstyle E},p_{\scriptscriptstyle E},\lambda^{2}_{a},\pi^{a})$, described by the action functional $\int dt(p_{i}\dot{q}^{i}+\pi^{a}\dot{\lambda}^{2}_{a}-H_{T}-\lambda^{\pi}_{a}\pi^{a})$. Effectively, this approach introduces conjugate momenta $\pi^{a}$ for the Lagrange multipliers $\lambda^{2}_{a}$ that are constrained to vanish on-shell. The total Hamiltonian of this system is given by $H_{T}+\lambda^{\pi}_{a}\pi^{a}$. Yet, the crucial difference here is that the standard velocity-momentum relations are, again, no longer satisfied. Instead, the equations of motion $\dot{q}^{\scriptscriptstyle E}_{A}=\{q^{\scriptscriptstyle E}_{A},H_{E}\}$, together with the relation $f_{A}^{-1}(q^{\scriptscriptstyle E},p_{\scriptscriptstyle E}^{A},\dot{q}^{\scriptscriptstyle E}_{a})=\frac{\partial H_{can}}{\partial p_{A}}(q^{\scriptscriptstyle E},p^{A}_{\scriptscriptstyle E})$, lead to $p^{A}_{\scriptscriptstyle E}=w^{A}(q^{\scriptscriptstyle E},\dot{q}^{\scriptscriptstyle E},\lambda^{2}_{a})$, where $w^{A}$ is the solution to the equations
$\dot{q}^{\scriptscriptstyle E}_{A}-f^{-1}_{A}(q^{\scriptscriptstyle E},w^{B}(q^{\scriptscriptstyle E},\dot{q}^{\scriptscriptstyle E},\lambda^{2}_{a}))+\lambda^{2}_{a}\frac{\partial\phi^{a}_{1}}{\partial p_{\scriptscriptstyle E}^{A}}(q^{\scriptscriptstyle E},w^{B}(q^{\scriptscriptstyle E},\dot{q}^{\scriptscriptstyle E},\lambda^{2}_{a}))=0.$ (76)
Consequently, the standard Lagrangian is replaced by the following expression:
$L_{E}=L(q^{\scriptscriptstyle E},f^{-1}_{A}(q^{\scriptscriptstyle E},w^{B},\dot{q}^{\scriptscriptstyle E}_{a}),\dot{q}^{\scriptscriptstyle E}_{a})+w^{A}(q^{\scriptscriptstyle E},\dot{q}^{\scriptscriptstyle E},\lambda^{2}_{a})\left(\dot{q}^{\scriptscriptstyle E}_{A}-f^{-1}_{A}(q^{\scriptscriptstyle E},w^{B})\right)-\lambda^{2}_{a}\phi^{a}_{1}(q^{\scriptscriptstyle E},w^{B}).$ (77)
The Euler–Lagrange equations derived from $L_{E}$ are fully equivalent to the equations of the extended Hamiltonian system. This explicitly demonstrates that the dynamics of the original variables are altered – a conclusion that remains valid for arbitrary $\lambda^{2}_{a}$, not only for small perturbations. A detailed treatment of this result is provided in [25], where $L_{E}$ is termed the extended Lagrangian. In the same work, this Lagrangian is employed as a tool for proving Dirac’s conjecture: it is established that all first-class constraints independently generate local symmetry transformations of the extended Lagrangian system.
@@
This discussion generalizes to situations where the gauge “hits” arbitrarily many times. Specifically, the tertiary constraints $\phi^{a}_{3}$ may arise from the preservation of the secondary constraints $\phi^{a}_{2}$, followed by further constraints originating from the consistency conditions of the tertiary ones, and so on, with all the constraints in the resulting chain being first-class [^22] An example of a system with tertiary constraints $L=\dot{x}\dot{z}+yz$ can be found in [19].. The reasoning in such cases is analogous: one performs a procedure similar to (72) that incorporates all first-class constraints in the chain. For example, for a system whose chain ends at the tertiary level, we have:
$q_{i}=q^{\scriptscriptstyle E}_{i}-\lambda^{2}_{a}\{q^{\scriptscriptstyle E}_{i},\phi_{1}^{a}\}-\lambda^{3}_{a}\{q^{\scriptscriptstyle E}_{i},\phi^{a}_{2}\},\qquad p^{i}=p_{\scriptscriptstyle E}^{i}-\lambda^{2}_{a}\{p_{\scriptscriptstyle E}^{i},\phi_{1}^{a}\}-\lambda^{3}_{a}\{p_{\scriptscriptstyle E}^{i},\phi^{a}_{2}\}.$ (78)
@@
The tertiary constraints can be chosen as $\phi^{a}_{3}\equiv\{\phi^{a}_{2},H_{can}\}$. Thus, after performing this substitution and retaining terms only to first order in $\lambda^{2}_{a}$ and $\lambda^{3}_{a}$, one again arrives at $L^{(1)}=p_{\scriptscriptstyle E}^{i}\dot{q}^{\scriptscriptstyle E}_{i}-H_{E}$.
@@
Moreover, the above results extend to systems that possess first-class constraints together with second-class ones. Second–class constraints make it possible to impose restrictions on variables belonging to a single canonically conjugate pair simultaneously and to eliminate these variables from the equations. This elimination can be carried out simply by direct substitution into the total Hamiltonian [6], thereby reducing the system to one with only first-class constraints [^23] Second–class constraints also correspond to constrained modes. Eliminating the constraints in this manner does not remove the corresponding information they carry from the theory. They must still be accounted for, since they often correspond to observable quantities [17].. In practice, however, this procedure may turn out to be rather cumbersome.
@@
Thus, we have shown that all our arguments apply to all canonical mechanical gauge theories [^24] A trivial exception occurs in theories in which the gauge “hits” only once. The only possible first-class constraints in such theories are primary ones. In that case the total and extended Hamiltonians coincide, and Dirac’s conjecture is automatically satisfied in both versions. An example of this is a relativistic free particle.. Carrying out an analogous generalization for well-behaved field theories also presents no difficulty.
@@ 6 Conclusions
In this work, we have explored certain aspects of the Dirac–Bergmann approach to the Hamiltonian formulation of gauge theories that, in our view, do not receive the attention they deserve in the literature. We conclude that the question of gauge symmetry generators for degenerate Hamiltonian systems, and the treatment of such systems within the extended Hamiltonian framework, are more subtle than they may appear. Ignoring these nuances leads to a naïve interpretation in which the canonical variables are assumed to play the same role as in the total Hamiltonian approach. In this case, the system appears to possess twice the actual gauge symmetry, since Dirac’s conjecture is fulfilled, with the constrained modes becoming spurious.
Although there is nothing inherently wrong with introducing additional gauge freedom into a theory, and even though such a naïve interpretation leaves the dynamical modes intact, it is essential to account for the constrained part of the physical sector, as it corresponds to non-dynamical observables. Constrained modes must therefore be expressed in terms of combinations that commute with the first -class constraints in the extended Hamiltonian approach. This actually implies that we must redefine the system’s observables. Interestingly, this redefinition can be interpreted as a version of the Stückelberg trick applied to variables that are second-class with respect to the set of all primary first-class constraints.